# Online Convex Optimization with Memory and Limited Predictions

### v1

@@ 1 Pick $u^{0}\sim\widetilde{\mathcal{N}}(0,I_{d})$
3       Let $r=t+W-1$.
4       Set $\bar{x}_{r}^{0}=x_{r}^{0}+\delta u^{0}$ and $\tilde{x}_{r}^{0}=x_{r}^{0}+\delta u^{0}$.
@@
7       Update $x_{r+1}^{0}=\mathbf{\Pi}_{\mathcal{X}}(x_{r}^{0}-\eta_{r}g_{r}^{0})$.
@@
13                   Set $\bar{x}_{s+h-1}^{j}=x_{s+h-1}^{j}+\delta^{\prime}u_{s}^{j}$ and $\tilde{x}_{s+h-1}^{j}=x_{s+h-1}^{j}+\delta^{\prime}u_{s}^{j}$.
@@
18            Update $x_{s}^{j+1}=\mathbf{\Pi}_{\mathcal{X}}\big{(}x_{s}^{j}-\alpha\sum_{k=s}^{s+h-1}g_{k}^{j}\big{)}$.
@@
20             Set $\bar{x}_{s}^{j+1}=x_{s}^{j+1}+\delta^{\prime}u_{s}^{j+1}$ and $\tilde{x}_{s}^{j+1}=x_{s}^{j+1}+\delta^{\prime}u_{s}^{j+1}$.
@@
One can also check that all the function values in $\mathcal{I}_{t}$ are available via the predictions in lines 5, 14 and 20 of Algorithm 1. Moreover, at any $t\in\{2-W,\dots,T\}$, Algorithm 1 only needs to maintain in its memory those functions values in $\mathcal{I}_{t}$ (that are needed to obtain $x_{t}^{K}$) and remove the other function values from its memory. In other words, the memory of Algorithm 1 may be recursively updated through receiving fresh predictions via $l_{t}(\cdot)$ and removing the function values of $l_{t}(\cdot)$ that are no longer useful. Finally, we note that for $t=2-W,\dots,T$, the points associated with the predictions in lines 5, 14 and 20 of Algorithm 1 correspond to consecutive points on $2(K+1)$ trajectories given by $\{\bar{x}_{1}^{j},\bar{x}_{2}^{j},\dots\}_{j=0}^{K}$ and $\{\tilde{x}_{1}^{j},\tilde{x}_{2}^{j},\dots\}_{j=0}^{K}$, and thus we do not need to restart the simulator $l_{t}(\cdot)$ during the course of Algorithm 1.
Truncated Gaussian Smoothing. In Algorithm 1 (lines 4, 13 and 19), we use a truncated Gaussian random vector denoted as $u_{s}^{j}\sim\widetilde{\mathcal{N}}(0,I_{d})$ to perturb the point $x_{s}^{j}$ to obtain $\bar{x}_{s}^{j}$ and $\tilde{x}_{s}^{j}$ which are then fed to the simulator $l_{t}(\cdot)$. Specifically, for any $i\in[d]$, we let the $i$th margin $u_{s,i}^{j}$ of $u_{s}^{j}$ to be a truncated Gaussian random variable obtained from the normal distribution such that $u_{s,i}^{j}$ belongs to the interval $[-(1/2d^{2}(2h-1))^{1/4},1/(2d^{2}(2h-1))^{1/4}]$ and the pdf of $u_{s,i}^{j}$ (i.e., the $i$th margin of $\widetilde{\mathcal{N}}(0,I_{d})$) is given by
@@
It follows that $\lVert u_{s}^{j}\rVert\leq 1/(2(2h-1))^{1/4}$. Formally, we introduce the following [39].
@@
There exist efficient methods for sampling from a truncated Gaussian distribution (see e.g. [40]). As we discussed in the introduction, using the truncated Gaussian allows us to prove the desired regret guarantee of Algorithm 1 in the next section.
@@ 4 Regret Analysis
We now analyze the regret of Algorithm 1. To proceed, we introduce the following definitions and assumptions, which are standard in the OCO literature [6, 15, 8, 5].
@@
$\left\|\nabla f(x)-\nabla f(y)\right\|\leq\beta\left\|x-y\right\|,\ \forall x,y\in\mathbb{R}^{d}.\$
@@
The feasible set $\mathcal{X}\subseteq\mathbb{R}^{d}$ is convex and satisfies that $\max_{x,y\in\mathcal{X}}\left\|x-y\right\|\leq D$.
@@
Note that the noisy model in Assumption 3 is equivalent to assuming that $l_{t}(\bm{x})=f_{t}(\bm{x})+\phi_{t}$ for all $\bm{x}\in\mathcal{X}^{h}$ with $\phi_{t}\in\mathbb{R}$ and $\left|\phi_{t}\right|\leq\varphi_{t}$, where $\phi_{t}$ can be either deterministic or stochastic. The bounded noise model has also been considered in e.g. [33]. The potentially unbounded stochastic noise model has been considered in e.g. [41] and we leave the extension to such a noise model to future work.
@@
The following result shows that the functions $\hat{f}_{t}^{c}(\cdot)$ and $\hat{C}_{T}(\cdot)$ inherit the properties from $f_{t}(\cdot)$; the proof can be found in Appendix A.
@@
Suppose Assumption 1(a) holds for $f_{t}(\cdot)$ for all $t\in[T]$. Then, (a) $\hat{f}_{t}^{c}(\cdot)$ is $\beta$-smooth and $\mu$-strongly convex; (b) $C_{T}(\cdot)$ is $\beta h$-smooth, $\mu$-strongly convex and $G\sqrt{Th}$-Lipschitz; and (c) $\hat{C}_{T}(\cdot)$ is $\beta h$-smooth and $\mu$-strongly convex.
@@
Under Assumption 1(a), the following hold:
(a) For $f_{t}^{c}(\cdot)$ and $\hat{f}_{t}^{c}(\cdot)$ given by Definition 5(a), any $x\in\mathbb{R}^{d}$ and any $\delta\in\mathbb{R}_{>0}$,
@@
$\frac{\partial\hat{C}_{T}(\bm{x})}{\partial x_{s}}=\mathbb{E}_{v_{s}}\Big{[}\sum_{k=s}^{s+h-1}\frac{f_{k}(\bar{x}_{k-h+1:k})-f_{k}(\tilde{x}_{k-h+1:k})}{2\delta^{\prime}}v_{s}\Big{]}\ \forall s\in[T],$ (8)
@@
Initialization Phase: We first upper bound the contribution of the initialization step in Algorithm 1 to $R(\mathcal{A})$, where we note from the definition of Algorithm 1 that $x_{1}^{0},\dots,x_{T}^{0}$ are chosen by the initializing phase of the algorithm. The proof of Theorem 1 can be found in Appendix B.
@@
Suppose Assumptions 1-3 hold. Let $\eta_{t}=\frac{1}{t\mu}$ and $\delta=1/\sqrt{T}$ in Algorithm 1. Then, the initialization step of Algorithm 1 satisfies that
$\mathbb{E}\Big{[}\sum_{t=1}^{T}f_{t}(x_{t-h+1:t}^{0})-f_{t}(x_{t-h+1:t}^{\star})\Big{]}\leq\Big{(}\frac{\sqrt{2}D}{\mu}+Gh^{2}\Big{)}V_{T}+T\delta^{2}\beta d$
$+(\frac{8G^{2}h^{2}+hG^{2}}{2\mu(2h-1)}+\frac{h^{3}G^{2}D}{\delta(2(2h-1))^{1/2}}+\frac{3G^{2}h^{3}}{\mu(2(2h-1))^{1/2}})(1+\ln T)$
$+\frac{2}{\delta^{2}\mu\sqrt{2h-1}}\sum_{t=1}^{T}\varphi_{t}^{2}+(\frac{\sqrt{2}h^{2}GD}{2\delta(2(2h-1))^{1/2}}+\frac{\sqrt{2}Gh^{2}+D\mu}{\delta\mu(2(2h-1))^{1/4}})\sum_{t=1}^{T}\varphi_{t}$ (11)
with $V_{T}\triangleq\mathbb{E}\big{[}\sum_{t=1}^{T-1}\left\|x_{t}^{\star}-x_{t+1}^{\star}\right\|\big{]}$, where $\mathbb{E}[\cdot]$ denotes the expectation with respect to the randomness in Algorithm 1 and $(x_{1}^{\star},\dots,x_{T}^{\star})$ is an optimal solution to (1).
@@
$x_{t+1}^{0}=\Pi_{\mathcal{X}}(x_{t}^{0}-\eta_{t}g_{t}^{0}),$
where $g_{t}^{0}$ is given by line 6 of Algorithm 1. As we argued in Remark 1, our OCO with memory framework corresponds to the bandit information setting, i.e., $g_{t}^{0}$ is computed based on only the (noisy) cost function values. The resulting dynamic regret of the initialization step applied to the problem of bandit OCO with memory is then given by (11) in Theorem 1. The OCO with memory framework has been studied in [12, 13, 8]. We extend the existing algorithm design and regret analysis to the bandit OCO with memory framework, where the function value oracle can also be noisy. In particular, setting $\varphi_{t}=0$ for all $t\geq 0$ (i.e., there is no noise in the oracle of the function values of $f_{t}(\cdot)$), the regret bound in (11) reduces to $\tilde{\mathcal{O}}(V_{T}\sqrt{T})$ with $\tilde{\mathcal{O}}(\cdot)$ hiding logarithmic factors in $T$, where the scaling of $T$ matches the regret bound for the algorithm proposed in [8] for (full information) OCO with memory (i.e., the cost functions are fully accessible). Note that the $V_{t}$ factor in the regret bound in (11) can potentially be refined to $\sqrt{V_{T}}$ using the techniques from [10, 8]. Specifically, supposing $V_{T}$ (or an upper bound on $V_{T}$) is known, then setting the step size $\eta=\frac{1}{t\mu\sqrt{V_{T}}}$ in Algorithm 1 yields a regret bound of $\tilde{\mathcal{O}}(\sqrt{V_{T}T})$.
Zeroth-Order Update Phase: Next, we upper bound the contribution of the zeroth-order method used in Algorithm 1 to $R(\mathcal{A})$. As we argued before in Section 3, the iterates $\bm{x}^{j}=(x_{1}^{j},\dots,x_{T}^{j})$ for $j=0,\dots,K-1$ in Algorithm 1 are obtained from the update rule $\bm{x}^{j+1}=\bm{x}^{j}-\alpha\bm{g}^{j}$ starting from the initial point $\bm{x}^{0}$ (given by the initialization phase of Algorithm 1), where $\bm{g}^{j}=(\bm{g}^{j}_{1},\dots,\bm{g}^{j}_{T})\in\mathbb{R}^{dT}$ with $\bm{g}^{j}_{s}=\sum_{k=s}^{s+h-1}g_{k}^{j}$ for all $s\in[T]$ and $\bm{g}^{j}$ can be viewed as an estimate of the true gradient $\nabla C_{T}(\bm{x}^{j})$ based on the function values of $l_{t}(\cdot)$ evaluated at certain points perturbed by the truncated Gaussian smoothing (see our arguments in Section 3). Thus, the following result indeed characterizes the convergence of the above update rule to the optimal solution to problem (1), given the initial point $\bm{x}^{0}=(x_{1}^{0},\dots,x_{T}^{0})$ returned by the initialization phase of Algorithm 1. The proof of Theorem 2 can be found in Appendix B.
@@
Consider the same setting as Theorem 1 and further let $\alpha=1/\beta^{\prime}$ in Algorithm 1, where $\beta^{\prime}=\beta h$ with $\beta$ given by Assumption 1. The output $\bm{x}^{K}=(x_{1}^{K},\dots,x_{T}^{K})$ of Algorithm 1 satisfies that
@@
$\mathbb{E}\big{[}C_{T}(\bm{x}^{K})-C_{T}(\bm{x}^{\star})\big{]}\leq\Big{(}\frac{1}{1+\gamma}\Big{)}^{K}\big{(}C_{T}(\bm{x}^{0})-C_{T}(\bm{x}^{\star})\big{)}+\frac{\varepsilon}{\gamma},$ (12)
@@
where $\gamma\triangleq\mu/(\beta^{\prime}-\mu)$ [^3] For any $\beta$-smooth and $\mu$-strongly convex function $f(\cdot)$, it is known that $\beta>\mu$ holds (e.g., [30]). and
@@
$\varepsilon\triangleq\frac{D\delta^{\prime}\beta h\sqrt{T}}{2\sqrt{2}(2(2h-1))^{3/4}}+\sqrt{h\beta GT\delta^{\prime}}D(Td+3)^{3/4}+\frac{Dh\sum_{t=1}^{T}\varphi_{t}}{\delta^{\prime}(2(2h-1))^{1/4}}.$

### v3

@@ 1 Pick $u^{0}\sim\widetilde{\mathcal{N}}(0,I_{d})$
3   Let r=t+W1r=t+W-1.
4   Set $\bar{x}_{r}^{0}=x_{r}^{0}+\delta u^{0}$ and $\tilde{x}_{r}^{0}=x_{r}^{0}-\delta u^{0}$.
@@
7   Update $x_{r+1}^{0}=\mathbf{\Pi}_{(1-\xi)\mathcal{X}}(x_{r}^{0}-\eta_{r}g_{r}^{0})$.
@@
13         Set $\bar{x}_{s+h-1}^{j}=x_{s+h-1}^{j}+\delta^{\prime}u_{s}^{j}$ and $\tilde{x}_{s+h-1}^{j}=x_{s+h-1}^{j}-\delta^{\prime}u_{s}^{j}$.
@@
18     Update $x_{s}^{j+1}=\mathbf{\Pi}_{(1-\xi^{\prime})\mathcal{X}}\big(x_{s}^{j}-\alpha\sum_{k=s}^{s+h-1}g_{k}^{j}\big)$.
@@
20      Set $\bar{x}_{s}^{j+1}=x_{s}^{j+1}+\delta^{\prime}u_{s}^{j+1}$ and $\tilde{x}_{s}^{j+1}=x_{s}^{j+1}-\delta^{\prime}u_{s}^{j+1}$.
@@
Time Complexity. The time complexity of Algorithm 1 at each time step $t\geq 1$ is dominated by the second subroutine, which performs $K=\lfloor\frac{W}{h-1}\rfloor$ update steps (line 17). Since these updates are derivative-free, the complexity is primarily determined by the number of function value queries needed for predictions. We will see in Section 4 that these update steps enable the exponential decay in regret with the prediction window size $W$. Finally, the entire algorithm operates without gradient feedback because it is fundamentally designed for the bandit feedback setting. A core premise of our problem is that the decision-maker cannot access the analytical form of the cost functions or their gradients. Interaction is limited to querying the system with an action and observing the resulting cost value. To adhere to this practical yet challenging information constraint, both subroutines are necessarily designed as derivative-free methods.
@@
One can also check that all the function values in $\mathcal{I}_{t}$ are available via the predictions in lines 5, 14 and 20 of Algorithm 1. Moreover, at any $t\in\{2-W,\dots,T\}$, Algorithm 1 only needs to maintain in its memory those function values in $\mathcal{I}_{t}$ (that are needed to obtain $x_{t}^{K}$) and remove the other function values from its memory. In other words, the memory of Algorithm 1 may be recursively updated through receiving fresh predictions via $l_{t}(\cdot)$ and removing the function values of $l_{t}(\cdot)$ that are no longer useful. Finally, we note that for $t=2-W,\dots,T$, the points associated with the predictions in lines 5, 14 and 20 of Algorithm 1 correspond to consecutive points on $2(K+1)$ trajectories given by $\{\bar{x}_{1}^{j},\bar{x}_{2}^{j},\dots\}_{j=0}^{K}$ and $\{\tilde{x}_{1}^{j},\tilde{x}_{2}^{j},\dots\}_{j=0}^{K}$, and thus we do not need to restart the simulator $l_{t}(\cdot)$ during the course of Algorithm 1.
Truncated Gaussian Smoothing. In Algorithm 1 (lines 4, 13 and 19), we use a truncated Gaussian random vector denoted as $u_{s}^{j}\sim\widetilde{\mathcal{N}}(0,I_{d})$ to perturb the point $x_{s}^{j}$ to obtain $\bar{x}_{s}^{j}$ and $\tilde{x}_{s}^{j}$ which are then fed to the simulator $l_{t}(\cdot)$. Specifically, for any $i\in[d]$, we let the $i$th margin $u_{s,i}^{j}$ of $u_{s}^{j}$ to be a truncated Gaussian random variable obtained from the normal distribution such that $u_{s,i}^{j}$ belongs to the interval $(1/2d^{2}(2h-1))^{1/4},1/(2d^{2}(2h-1))^{1/4}]$ and the probability density function (pdf) of $u_{s,i}^{j}$ (i.e., the $i$th margin of $\widetilde{\mathcal{N}}(0,I_{d})$) is given by
@@
It follows that $\lVert u_{s}^{j}\rVert\leq 1/(2(2h-1))^{1/4}\leq 1$, since $h>1$. Formally, we introduce the following johnson1995continuous.
@@
There exist efficient methods for sampling from a truncated Gaussian distribution (see e.g. robert1995simulation). As we discussed in the introduction, using the truncated Gaussian allows us to prove the desired regret guarantee of Algorithm 1 in the next section.
@@ 4 Regret Analysis
We now analyze the regret of Algorithm 1. To proceed, we introduce the following definitions and assumptions, which are standard in the OCO literature agarwal2010optimal; agarwal2019online; zhao2022non; flaxman2005online.
@@
$\left\|\nabla f(x)-\nabla f(y)\right\|\leq\beta\left\|x-y\right\|,\ \forall x,y\in\mathbb{R}^{d}.\\$
@@
The feasible set $\mathcal{X}\subseteq\mathbb{R}^{d}$ is convex, contains the origin, and satisfies $\max_{x,y\in\mathcal{X}}\left\|x-y\right\|\leq D$.
Remark 1.
As we discussed in Section 3, Algorithm 1 (lines 7 and 17) projects onto smaller sets $(1-\xi)\mathcal{X}$ and $(1-\xi^{\prime})\mathcal{X}$, where $\xi,\xi^{\prime}\in\mathbb{R}_{\geq 0}$. By Assumption 2, we know that $(1-\xi)\mathcal{X}$ is also convex and contains the origin. Moreover, for any two $y,z\in(1-\xi)\mathcal{X}$, we can write $y=(1-\xi)x_{1}$ and $z=(1-\xi)x_{2}$ for some $x_{1},x_{2}\in\mathcal{X}$, which implies via Assumption 2 that $\left\|y-z\right\|=\left\|(1-\xi)(x_{1}-x_{2})\right\|=(1-\xi)\left\|x_{1}-x_{2}\right\|\leq(1-\xi)D$. Similar arguments hold for the set $(1-\xi^{\prime})\mathcal{X}$. Now, we formally show how to ensure that perturbed used points in lines 4,13,19 of Algorithm 1 stay within the set $\mathcal{X}$ as promised in Section 3. To this end, consider a perturbed point, e.g., $\bar{x}=x+\delta u$, where $x$ is kept within the smaller set $(1-\xi)\mathcal{X}$ via projection and $u\sim\tilde{\mathcal{N}}(0,I_{d})$ as per Definition 1. Recall that $\mathcal{X}$ contains the origin by Assumption 2, we know that there exists a ball $r\mathbb{B}=\{x:\left\|x\right\|\leq r\}$ with radius $r>0$ such that $r\mathbb{B}\subseteq\mathcal{X}$, which implies that any point $y\in(1-\xi)\mathcal{X}$ has a distance of at least $\xi r$ to the boundary of $\mathcal{X}$. Recalling that $\left\|u\right\|\leq 1$ by definition, to guarantee that $\bar{x}=x+\delta u$ remains within $\mathcal{X}$, we can choose $\delta\leq\xi r$. Finally, it is worth mentioning that in the two phases of Algorithm 1, we use potentially different $\xi,\xi^{\prime}$ to obtain the sets $(1-\xi)\mathcal{X},(1-\xi^{\prime})\mathcal{X}$, and potentially different $\delta,\delta^{\prime}$ to obtain the perturbed points. Such choices of $\xi,\xi^{\prime},\delta,\delta^{\prime}$ facilitate the subsequent proof of the algorithm’s regret upper bounds.
@@
Note that the noisy model in Assumption 3 is equivalent to assuming that $l_{t}(\bm{x})=f_{t}(\bm{x})+\phi_{t}$ for all $\bm{x}\in\mathcal{X}^{h}$ with $\phi_{t}\in\mathbb{R}$ and $\left|\phi_{t}\right|\leq\varphi_{t}$, where $\phi_{t}$ can be either deterministic or stochastic. The bounded noise model has also been considered in e.g. berahas2022theoretical. The potentially unbounded stochastic noise model has been considered in e.g. agarwal2011stochastic, and we leave the extension to such a noise model to future work.
@@
The following result shows that the functions $\hat{f}_{t}^{c}(\cdot)$ and $\hat{C}_{T}(\cdot)$ inherit the properties from $f_{t}(\cdot)$; the proof is included in Appendix A.
@@
Suppose Assumption 1 holds for $f_{t}(\cdot)$ for all $t\in[T]$. Then, (a) $\hat{f}_{t}^{c}(\cdot)$ is $\beta$-smooth and $\mu$-strongly convex; (b) $C_{T}(\cdot)$ is $\beta h$-smooth, $\mu$-strongly convex and $G\sqrt{Th}$-Lipschitz; and (c) $\hat{C}_{T}(\cdot)$ is $\beta h$-smooth and $\mu$-strongly convex.
@@
Under Assumption 1(a), the following hold: (a) For $f_{t}^{c}(\cdot)$ and $\hat{f}_{t}^{c}(\cdot)$ given by Definition 5(a), any $x\in\mathbb{R}^{d}$ and any $\delta\in\mathbb{R}_{>0}$,
@@
$\frac{\partial\hat{C}_{T}(\bm{x})}{\partial x_{s}}=$ $\mathbb{E}_{v_{s}}\Big[\sum_{k=s}^{s+h-1}\frac{f_{k}(\bar{x}_{k-h+1:k})-f_{k}(\tilde{x}_{k-h+1:k})}{2\delta^{\prime}}v_{s}\Big]\qquad\forall s\in[T],$ (8)
@@
Initialization Phase Regret. We first upper bound the contribution of the initialization step in Algorithm 1 to $R(\mathcal{A})$, where we note from the definition of Algorithm 1 that $x_{1}^{0},\dots,x_{T}^{0}$ are chosen by the initializing phase of the algorithm.
Note that the modification of projecting onto a smaller set $(1-\xi)\mathcal{X}$ ensures the feasibility of query points in Remark 1. To analyze the regret while avoiding dependencies on the variation (also called path length) of the specific optimal solution within the smaller set, we utilize a scaled version of the true optimal sequence as our comparator. Let $\bm{x}^{\star}=(x_{1}^{\star},\dots,x_{T}^{\star})$ be the optimal solution to (1). We define a scaled comparator sequence $\hat{\bm{x}}^{\star}=(\hat{x}_{1}^{\star},\dots,\hat{x}_{T}^{\star})$ where $\hat{x}_{t}^{\star}=(1-\xi)x_{t}^{\star}$. Since $\mathcal{X}$ is convex and contains the origin, $\hat{\bm{x}}^{\star}$ lies within the smaller feasible set $((1-\xi)\mathcal{X})^{T}$ as discussed in Remark 1. We then decompose the initialization regret as follows:
$R(\mathcal{A}_{\mathrm{init}})$ $=C_{T}(\bm{x}^{0})-C_{T}(\bm{x}^{\star})$
$=\underbrace{\left(C_{T}(\bm{x}^{0})-C_{T}(\hat{\bm{x}}^{\star})\right)}_{\text{Regret against scaled comparator}}+\underbrace{\left(C_{T}(\hat{\bm{x}}^{\star})-C_{T}(\bm{x}^{\star})\right)}_{\text{Approximation Error}}.$
The following theorem bounds the regret in the initialization phase for Algorithm 1, utilizing the fact that the path length of the scaled comparator is bounded by the path length of the true optimum. The detailed proof can be found in Appendix B.
@@
Suppose Assumptions 1-3 hold. Let $\eta_{t}=\frac{1}{t\mu}$ in Algorithm 1. Then, the initialization step of Algorithm 1 satisfies that
$\mathbb{E}[R(\mathcal{A}_{\text{init}})]$ $\leq\Big(\frac{\sqrt{2}(1-\xi)^{2}D}{\mu}+(1-\xi)Gh^{2}\Big)V_{T}+T\delta^{2}\beta d$
$\quad+\Big(\frac{8G^{2}h^{2}+hG^{2}}{2\mu(2h-1)}+\frac{h^{3}G^{2}(1-\xi)D}{\delta(2(2h-1))^{1/2}}+\frac{3G^{2}h^{3}}{\mu(2(2h-1))^{1/2}}\Big)(1+\ln T)$
$\quad+\frac{2}{\delta^{2}\mu\sqrt{2h-1}}\sum_{t=1}^{T}\varphi_{t}^{2}+\Big(\frac{\sqrt{2}h^{2}G(1-\xi)D}{2\delta(2(2h-1))^{1/2}}+\frac{\sqrt{2}Gh^{2}+(1-\xi)D\mu}{\delta\mu(2(2h-1))^{1/4}}\Big)\sum_{t=1}^{T}\varphi_{t}$
$\quad+\xi GT\sqrt{h}D,$ (11)
with $V_{T}\triangleq\mathbb{E}\big[\sum_{t=1}^{T-1}\left\|x_{t}^{\star}-x_{t+1}^{\star}\right\|\big]$, where $\mathbb{E}[\cdot]$ denotes the expectation with respect to the randomness in Algorithm 1 and $\bm{x}^{\star}$ is the global optimal solution defined in (2).
@@
$x_{t+1}^{0}=\Pi_{(1-\xi)\mathcal{X}}(x_{t}^{0}-\eta_{t}g_{t}^{0}),$
[…]
Zeroth-Order Update Phase. Next, we upper bound the contribution of the zeroth-order method used in Algorithm 1 to $R(\mathcal{A})$. As we argued before in Section 3, the iterates $\bm{x}^{j}=(x_{1}^{j},\dots,x_{T}^{j})$ for $j=0,\dots,K-1$ in Algorithm 1 are obtained from the update rule $\bm{x}^{j+1}=\mathbf{\Pi}_{(1-\xi^{\prime})\mathcal{X}}(\bm{x}^{j}-\alpha\bm{g}^{j})$ starting from the initial point $\bm{x}^{0}$ (given by the initialization phase of Algorithm 1), where $\bm{g}^{j}=(\bm{g}^{j}_{1},\dots,\bm{g}^{j}_{T})\in\mathbb{R}^{dT}$ with $\bm{g}^{j}_{s}=\sum_{k=s}^{s+h-1}g_{k}^{j}$ for all $s\in[T]$ and $\bm{g}^{j}$ can be viewed as an estimate of the true gradient $\nabla C_{T}(\bm{x}^{j})$ based on the function values of $l_{t}(\cdot)$ evaluated at certain points perturbed by the truncated Gaussian smoothing (see our arguments in Section 3).
The following result characterizes the total dynamic regret of Algorithm 1 by combining the initialization regret and the convergence properties of the zeroth-order update. The detailed proof, including the decomposition of regret and the convergence analysis, is provided in Appendix B.
@@
Consider the same setting as Theorem 1 and further let $\alpha=1/\beta^{\prime}$ in Algorithm 1, where $\beta^{\prime}=\beta h$ with $\beta$ given by Assumption 1. Then, the overall regret $R(\mathcal{A})$ of Algorithm 1 defined in (3) satisfies:
@@
$\mathbb{E}[R(\mathcal{A})]\leq\Big(\frac{1}{1+\gamma}\Big)^{K}$ $\mathbb{E}[C_{T}(\bm{x}^{0})-C_{T}(\bm{x}^{\star})]+\frac{\varepsilon}{\gamma}+\xi^{\prime}GT\sqrt{h}D,$ (12)
@@
where $\mathbb{E}[C_{T}(\bm{x}^{0})-C_{T}(\bm{x}^{\star})]$ is the initialization regret given in Theorem 1, $\gamma\triangleq\mu/(\beta^{\prime}-\mu)$ [^3] For any $\beta$-smooth and $\mu$-strongly convex function $f(\cdot)$, it is known that $\beta>\mu$ holds (e.g., nesterov2013introductory)., and
@@
$\varepsilon\triangleq\frac{(1-\xi^{\prime})D\delta^{\prime}\beta h\sqrt{T}}{2\sqrt{2}(2(2h-1))^{3/4}}+\sqrt{h\beta GT\delta^{\prime}}(1-\xi^{\prime})D(Td+3)^{3/4}+\frac{(1-\xi^{\prime})Dh\sum_{t=1}^{T}\varphi_{t}}{\delta^{\prime}(2(2h-1))^{1/4}}.$
We see from Theorem 2 that for the zeroth-order update phase (lines 8-20) to be effective, the number of iterations $K$ must be at least $1$. Since $K=\lfloor W/(h-1)\rfloor$, a minimum requirement is that $W\geq h-1$. Note that we do not treat $W$ to be an input parameter to Algorithm 1; rather, the value of $W$ depends on the specific problem setup. In practice, $W$ captures the ability to obtain predictions using the simulator $l_{t}(\cdot)$. Specifically, Theorem 2 shows that the dynamic regret decays exponentially with $K$, and thus with $W$, which suggests a larger $W$ is always beneficial for achieving lower regret. However, in many real-world applications, obtaining reliable predictions becomes increasingly difficult as the prediction horizon grows, since the quality and availability of future information may degrade significantly for a longer prediction window.